OpenStudy (anonymous):

Fill in the next term in the sequence , , , , , 1/8 1/4 1/2 how would i find this oone

zepdrix (zepdrix):

So we have a Geometric Sequence. So we're multiplying by some value when we move from one term to the next. We call this our `common ratio` and denote it by r. We can divide the first term by the second term, and that will tell us how much we multiplied by to get from one to the next.\[\large\rm \frac{a_2}{a_1}=r\]

OpenStudy (anonymous):

thats too complicated for my brain ..

zepdrix (zepdrix):

So what do you get when you divide? \[\large\rm \frac{\left(\frac{1}{4}\right)}{\left(\frac{1}{8}\right)}=?\]

OpenStudy (anonymous):

2?

zepdrix (zepdrix):

Yes, so our common ratio is 2. So to get the next term, simply multiply 1/2 by 2! :)